OpenStudy (anonymous):

3/2 (x+8) = 6x solving Equations Containing Fractions

OpenStudy (anonymous):

lets go step by step

OpenStudy (anonymous):

the \(\frac{3}{2}\) out front is annoying we can get rid of it by multpying both sides the the reciprocal, which is \(\frac{2}{3}\) because \[\frac{3}{2}\times \frac{2}{3}=1\]

OpenStudy (anonymous):

so the first step will be to multiply both sides by \(\frac{2}{3}\) to get \[x+8=6x\times \frac{2}{3}\] or \[4+8=4x\]

OpenStudy (anonymous):

hahahahaha you said it was annoying

OpenStudy (anonymous):

the whole prob is annoying

OpenStudy (anonymous):

now you have \(x\)'s on both sides of the equal sign you want them on one side only so you can arrange this by subtracting \(x\) from both sides that gives \[8=3x\]

OpenStudy (anonymous):

oh damn i see i made a typo above it should be \[x+8=4x\]

OpenStudy (anonymous):

so far so good

OpenStudy (anonymous):

now your last job is to divide both sides by \(3\) to get \(x\) by itself \[8=3x\] \[\frac{8}{3}=x\]

OpenStudy (anonymous):

2.6666666

OpenStudy (anonymous):

that doesnt look right